OpenStudy (anonymous):

Jackie invests 25000 part at 5.5% annual interest and the balance at 8.3% annual interest. How much is invested at each rate if jackie receives a 1 year interest payment of 1571?

OpenStudy (ttp):

Two variables, lets call the first X and the other is obviously the remainder from 25000 so subtract X and its the second variable: 1 ~ (X) 2 ~ (25000 - X) So now we need to set up the interest equation like so: Interest Gained = Amount[1] * Interest Rate [1] + Amount[2] * Interest Rate[2] In this example the interest earned is given as 1571, we will call Amount[1] as the variable X with interest rate 8.3% and Amount[2] would be the remainder or 25000 - X at the interest rate of 5.5%. 1571 = (X)*0.083 + (25000 - X)*0.055 Now Just solve for X to get the amount invested at 8.3%: 1571 = 0.083X + 1375 - 0.055X 196 = 0.028X X = 7000 Solve equation 2 to get the amount invested at 5.5%: 1 ~ X = 7000 2 ~ 25000 - X = 25000 - 7000 = 18000 So he invested $7000 at 8.3% interest and $18000 at 5.5% interest. Did i help?

OpenStudy (anonymous):

yes thank you

OpenStudy (ttp):

Could i have a medal? &' your welcome